OpenStudy (anonymous):

What is 3 3/8 divided by 9? I solved it but did not get any of these a. 3/7 b. 4/8 c. 2/8 d. 3/8 I know it's simple math but I can't seem to figure it out.

OpenStudy (anonymous):

d. 3/8

OpenStudy (jdoe0001):

\(\bf 3\cfrac{3}{8} \div 9\quad ?\)

OpenStudy (anonymous):

yes. how would it be 3/8?

OpenStudy (jdoe0001):

\(\bf 3\cfrac{3}{8} \div 9\\\quad \\\quad \\ a \cfrac{b}{c} \implies \cfrac{a\times c + b}{c} \textit{ thus then}\\\quad \\\quad \\ 3\cfrac{3}{8} \implies \cfrac{3 \times 8 + 3}{8} \implies \cfrac{27}{8}\quad \\\quad \\ 3\cfrac{3}{8} \div 9 \implies \cfrac{27}{8} \div 9 \implies \cfrac{27}{8} \div \cfrac{9}{1} \implies \cfrac{27}{8} \times \cfrac{1}{9}\)

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

please vote mine as best response

OpenStudy (anonymous):

ooooh okay. thank you!

OpenStudy (jdoe0001):

yw